Circular No. 6568

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) M. J. Mumma, M. A. DiSanti, and N. Dello Russo, Goddard Space

Flight Center, NASA; K. Magee-Sauer, Rowan College of New Jersey; and M. Fomenkova, University of California at San Diego, report: “We detected parent volatiles in C/1995 O1 on 1996 Dec. 10–12 and 1997 Jan. 21.5 UT, using the NASA Infrared Telescope Facility (+ CSHELL cryogenic infrared spectrometer) on Mauna Kea. The species, the vibrational bands, and the detected lines are: CO [1–0, 10 lines, J′ = 0–7]; C_{2}H_{6} [ν_{7}, ^{r}Q_{0}- and ^{r}Q_{1}- branches]; five lines of H_{2}O were also detected. All lines are peaked on the nucleus, and all are extended; CO extends nearly to the ends of the eastwest slit, about 15′′ in either direction. CH_{4} ν_{3} R0 was searched for on Dec. 11.1 but was not detected. The following reductions refer to a box 1′′ × 1′′.4 centered on the nucleus. On Jan. 21.5, the rotational temperature for CO was ∼ 130 K, and the production rate was ∼ 3.2 × 10^{29} molecules/s. The peak CO line flux measured was 8.7 × 10^{−17} W m^{−2} in R4. A preliminary analysis of five lines of H_{2}O suggests a rotational temperature ∼ 40–100 K. For 70 K, we derive Q(H_{2}O) = 5 × 10^{30} molecules/s. The flux for the ^{r}Q_{0}- branch (2983.4 cm^{−1}) was 6.9 × 10^{−18} W m^{−2}, and that for the^{r}Q_{1}-branch was 8.8 × 10^{−18} W m^{−2}. Assuming fluorescence excitation of the^{r}Q_{0} upper levels from the K = 0 and K = 2 lower levels, the combined production rate for molecules in those levels is 2.8 × 10^{27} molecules/s. Fluorescence excitation of the upper levels of ^{r}Q_{1} leads to a combined production rate of 3.6 × 10^{27} molecules/s for molecules in the K = 1 and K = 3 lower levels. Their sum (6.4 × 10^{27} molecules/s) is a lower bound to the total production rate for C_{2}H_{6}. The mixing ratios are then H_{2}O:CO:C_{2}H_{6} = 100:6.4:0.13 on Jan. 21.5. If the K-ladders exhibit a 130-K distribution like CO, the total ethane-production rate could be larger by ∼ 3. On Dec. 11.1, the rotational temperature for CO was ∼ 70 K, and the production rate was 2 × 10^{29}; for C_{2}H_{6}, the production rate in K = 0–3 was 7.1 × 10^{27}.”

1997 CU_{26}

R. P. Binzel, Massachusetts Institute of Technology, reports that measurements of the visible spectrum (range 450-950 nm) of the probable Centaur 1997 CU_{26} (cf. MPEC 1997-D11), obtained on Feb. 25.2 UT using the 2.4-m Hiltner telescope of the MDM Observatory at Kitt Peak, show a spectral slope similar to that of an average D-type asteroid.

1997 February 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green